# A Survey on Unlearning in Large Language Models

### v1

@@ 3.2. Post-Training Unlearning
The main approach to unlearning is to modify the parameters of trained LLMs, which is commonly referred to as the “post-training” phase. This leads to two crucial questions: (1) How to modify the parameters? (2) Which parameters should be selected for modifying? For question (1), the modification of parameters in most methods is an optimization problem, so we will introduce their objective design in Section 3.2.1. While some methods adopt alternative strategies, we include important ones in Section 3.2.2. For question (2), the modified parameters can be all parameters of the model, part of the parameters, or newly introduced parameters. We will discuss how to select part of the parameters in Section 3.2.3 and how to incorporate new parameters in Section 3.2.4. Note that parameter strategies can be freely combined with the objectives in Section 3.2.1.
3.2.1. Objective Optimization
The core of an unlearning mechanism is the design of its objective for optimization, which dictates how the model’s parameters are adjusted to forget specific knowledge while minimizing the impact on general utility. Different from the unified next-token prediction loss during learning, the design of unlearning objective is quite more diverse. Based on the primary target of the designated objective, we can classify existing methods into three major categories: Text-based, Distribution-based, and Activation-based. Figure 4 compares the general pipeline of different objective categories.
@@ 3.2.1. Objective Optimization
Text-based: These kind of objectives are most intuitive, which aim at minimize or maximize the predicted likelihood of certain text. A representative baseline is Gradient Ascent (Jang et al., 2023a; Yao et al., 2024b). It reduces the prediction probabilities of forget set samples by directly negating their cross-entropy next-token prediction loss. NPO (Zhang et al., 2024b) employs preference-based objective, introduces a reference model to constrain parameter changes. Another approach (Maini et al., 2024; Mekala et al., 2025) focuses on improve likelihood of “substitute responses”, which reply to a query from the forget set appropriately, without disclosing targeted knowledge. Although simple and effective, these baselines often spill unlearning effect beyond forget set itself, which impairs the model’s overall performance. To alleviate this, WGA (Wang et al., 2024) and FPGA (Feng et al., 2024) introduce methods to apply different weights to various token positions within the sequences of forget set. More methods put attention on generate or select data. Some utilize a external LLM to generate substitute responses relevant to the query in forget set  (Mekala et al., 2025; Sinha et al., 2025; Xu et al., 2025). Patil et al. (2025) and Chang and Lee (2025) proposed methods for selecting core subsets from the original unlearning corpus.
@@
Distribution-based: Text-based targets must provide labels from limited vocabulary, which restricts the optimization space of the model during unlearning. In order to achieve more fine-grained unlearning, some methods aim to make the model’s output distribution converge to a reference distribution that aligns with the unlearning goals. The key of these objectives lies in how to construct reference distribution. ME (Yuan et al., 2025b) uses a uniform distribution over the entire vocabulary. For other methods, reference distribution can establish either by modifying data or manipulating logits. WHP (Eldan and Russinovich, 2024) substitutes the unlearning target with unrelated entities to get general knowledge distribution that does not contain target information. WPU (Liu et al., 2025a) improves upon WHP by incorporating diverse substitute entities, performing entity name restoration, and augmenting input prompts. On the other hand, RKLD (Wang et al., 2025a) taking the difference between the logits of a model finetuned on the unlearning set and that of original model as reference. Similar logits-aware approach is also adopted by Obliviate (Russinovich and Salem, 2025) and PerMU (Wang et al., 2025b). Distribution-based objectives generally shorten distance to target distribution via minimizing divergence. The most popular choice is KL divergence, but there are also methods using reverse KL (Wang et al., 2025a), JS divergence (Singh et al., 2025) or f-divergence (Wang et al., 2025f).
@@
Activation-based: Both text-based and distribution-based methods treat the entire model as a black box, computing losses at the output level and perform back propagation, which is inefficient in many cases. Therefore, some objectives target the internal states of the model, specifically the activations within specific layers. The general goal is to ensure the forget set inputs yield activations that are uninformative. RMU (Li et al., 2024) combines a forget loss that perturbs hidden activations of harmful data towards a fixed random direction. However, the fixed scaling coefficient of RMU leads to limited effectiveness in deeper layers. To overcome this issue,  Dang et al. (2025) introduces an adaptive scaling coefficient proportional to the $l_{2}$-norm of the original representation. Similarly, LUNAR (Shen et al., 2025) aims to redirect the forget data’s activations into a refusal region, so that the model consistently produces safe refusal responses. Guo et al. (2024) and Wang et al. (2025g) take advantage of mechanism interpretability by constructing the activation of the expected answer after unlearning, and reversely calculating the closed-form solution of parameter updates.
@@
Multi-Objective Combination: Beyond unlearning objective itself, most methods add an loss term of retain set in practice to avoid degradation of general performance (Maini et al., 2024; Yao et al., 2024b; Wang et al., 2023; Yuan et al., 2025b). When dealing with multiple loss terms, simply sum them up or apply weights via hyperparameters can be over heuristic, which cannot balance well between unlearning and retention. NGDiff (Jin et al., 2025) treats the combination between objectives as a multi-task optimization problem, achieving a better trade-off through precise normalization and dynamic learning rates. MOLLM (Pan et al., 2025) computes a common descent direction in the dual gradient space, yielding an update that simultaneously reduces influence of the target knowledge while preserving overall utility.
3.2.2. Others
A small portion of post-training unlearning methods do not update the parameters through optimization. We review several notable approaches here as a complementary perspective, broadly categorizing them into (1) parameter arithmetic operations and (2) SAE-based methods.
Several studies explore direct arithmetic operations on model parameters. Inspired by advances in task vectors for knowledge editing (Ilharco et al., 2022), some methods fine-tune an intermediate model and combine its parameters arithmetically with those of the original model. For instance, SKU (Liu et al., 2024d) fine-tunes a “bad model” to obtain a parameter deviation that opposes the unlearning objective. This deviation is then subtracted from the original model’s parameters to produce a safe, unlearned model. A similar strategy is adopted in Eldan and Russinovich (2024).
Since training a model of comparable size to the original is computationally intensive, two key refinements have been proposed. First, fine-tuning can be performed using parameter-efficient fine-tuning (PEFT) techniques, where unlearning is achieved by applying negation operations to relevant parameter-efficient modules (PEMs) (Zhang et al., 2023b). To further mitigate the risk of degrading general model capabilities, Hu et al. (2024) combine an “expert” PEM with an “anti-expert” PEM and derive a general capability vector for preservation. Second, as an alternative to fine-tuning, approximate negative models can be derived via subspace decomposition and projection techniques, such as the Gram–Schmidt orthogonalization used in UNLEARN (Lizzo and Heck, 2025) and the singular value decomposition (SVD) applied in Ethos (Gao et al., 2024).
Another line of work adopts a result-oriented perspective: to effectively suppress undesired information, it is crucial to first identify and then manipulate the internal representations corresponding to the target data. Several studies integrate sparse autoencoders (SAEs) (Ng, 2011) into specific model layers to enhance interpretability and isolate relevant features. For example, Farrell et al. (2024) identify features that strongly activate on the unlearn set while minimally impacting the retain set, and then clamp their activations to negative values during inference. Similarly, Wu et al. (2024) introduce a trainable codebook between the encoder and decoder of an SAE. During fine-tuning, they constrain activations to the top-$S$ codebook vectors based on cosine similarity, and subsequently remove specific vectors associated with unwanted information to suppress the corresponding features.
@@ 3.2.3. Localizing Parameters
Data based Gradient DEPN (Wu et al., 2023) $\nabla\mathcal{L}_{f}$
SSU (Dou et al., 2025) ++ gradient of random labeling loss.
Stoehr et al. (2024) ++ gradient of KL divergence of output on retain set before/after unlearning.
MemFlex (Tian et al., 2024) ++ cosine similarity of $\nabla\mathcal{L}_{f}$ and $\nabla\mathcal{L}_{r}$.
WAGLE (Jia et al., 2025) ++ element-wise product of $\nabla\mathcal{L}_{f}$ and $\nabla\mathcal{L}_{r}$.
KLUE (Yang et al., 2025) ++ superficial knowledge regularization.
Activation Selective Pruning (Pochinkov and Schoots, 2024) four statistics of activations when processing forget versus retain data.
REVS (Ashuach et al., 2025) Combination of activation strength and token association.
FALCON (Hu et al., 2025) Mutual information of activations of the unlearn and retain set.
Data free Heuristics RMU (Li et al., 2024) Experimental observation and hyperparameter search optimization.
Adaptive RMU (Dang et al., 2025)
Mechanism LUNAR (Shen et al., 2025) Knowledge storage mechanism (down-projection matrix of MLP layers) (Meng et al., 2022).
LaW (Wang et al., 2025g)
@@
Simply fine-tuning of all parameters in a model often leads to issues such as high computational costs and potential performance degradation (Dou et al., 2025). In several studies on interpretability and model editing (Meng et al., 2022; Geva et al., 2021), researchers have demonstrated that knowledge is associated with specific model weights, thus proposing methods to locate relevant parameters for more efficient updates. Various techniques, including causal tracing (Meng et al., 2022), attribution patching (Nanda, 2023), probing (Geva et al., 2023), and path patching (Goldowsky-Dill et al., 2023), have been directly applied in research on unlearning (Guo et al., 2025). Furthermore, depending on the specific unlearning scenarios and objectives, numerous studies have proposed different strategies for parameter localization. Based on whether task-specific data are required, we categorize these methods into two distinct classes, as summarized in Table 2.
@@
Data-based. There are various ways to select certain layers or neurons within the model. The most common selection criterion is the loss gradient w.r.t. the model parameters, calculated on the unlearn set ($D_{u}$), which is based on the intuition that parameters with larger gradient magnitudes are more influential and should be prioritized for updates. DEPN (Wu et al., 2023) calculates the cumulative gradient of the loss function of the unlearn set and selects the top-k neurons. As an updated work, SSU (Dou et al., 2025) adds a random labeling loss to define a composite loss function, which is a commonly used data augmentation to enhance the stability (Neelakantan et al., 2015). Meanwhile, several works consider the retain set when selecting parameters, reducing the impact of parameter updates on the retain set (Stoehr et al., 2024, 2024; Tian et al., 2024) (refer to Table 2 for details). Furthermore, Yang et al. (2025) point out that different questions may share the same answer and should avoid unconditionally unlearning the answer regardless of the context. Thus, they propose KLUE, which introduces a superficial knowledge regularization for accurate parameter localization.
@@
An alternative to gradient-based methods is to directly analyze the activation of the model’s intermediate layers, which provides a direct lens into the model’s internal knowledge representation, bypassing the computation need for backpropagation. The method of selective pruning (Pochinkov and Schoots, 2024) calculates an importance score for each neuron based on four statistics of its activations when processing unlearn versus retain data. In addition to the activation strength, REVS (Ashuach et al., 2025) also considers the rank of a target token when projecting the neuron to the vocabulary space by unembedding matrix. A lower rank value indicates a stronger association between the target token and the neuron. They show that the combination outperforms methods based solely on activations, token associations, and gradients. Another approach, FALCON (Hu et al., 2025), uses mutual information of activations of the unlearn and retain set, to identify layers where the hidden representations of forget and retain knowledge are least entangled, targeting these specific layers for modification.
@@
Data-free. Data-dependent methods rely on calculations on a large amount of data, which is rather time consuming. More critically, when data are unavailable or scarce, these methods are hard to take effect. Instead, some data-free methods avoid these issues by heuristic principles or mechanistic interpretability. Li et al. (2024) observe that it is sufficient to compute the loss only on layer $\ell$ and update gradients only on layers $\ell-2$, $\ell-1$ and $\ell$, and perform a hyperparameter search over the layer to select the best layer for fine-tuning. This setting is followed by Dang et al. (2025). Additionally, inspired by insights into knowledge storage mechanism of LLMs (Meng et al., 2022), LUNAR (Shen et al., 2025) and LaW (Wang et al., 2025g) select the down-projection matrix of the MLP layers to update. Heuristic approaches rely on simple, effective rules to select intervention sites, whereas mechanistic approaches target the specific internal circuits responsible for knowledge generation.
@@ 3.2.4. Incorporating New Structure
This type of method generally maintains the original ability of the model by freezing the existing parameters, and achieves forgetting by introducing new parameters or auxiliary structures. A straightforward idea is to insert a new module between two layers of the model and only fine-tune this module, including EUL (Chen and Yang, 2023) and GRUN (Ren et al., 2025), which is illustrated in Figure 5(a). This module has significantly fewer parameters compared to the original model, sometimes combined with structures like soft gate functions to improve performance (Ren et al., 2025). To deal with a sequence of unlearning requests, EUL and GRUN train a separate module on each unlearn task and design a fusion mechanism to merge all modules.
@@
More research focuses on Low-Rank Adaptation (LoRA) (Hu et al., 2022) and other parameter-efficient fine-tuning (PEFT) techniques, which adds parameter-efficient modules (PEMs) to the model (Figure 5(b)). However, standard LoRA lacks sufficient plasticity and often performs poorly in selective unlearning scenarios (Cha et al., 2024), which is followed by several key enhancements. Cha et al. (2024) introduce Fisher-weighted Initialization of Low-rank Adapters (FILA). Meanwhile, Gao et al. (2025) address the challenge of continuous unlearning requests in practical settings. They employ an orthogonal regularization loss to disentangle different unlearning tasks within a single LoRA adapter and additionally train an out-of-distribution (OOD) detector to modulate the adapter activation based on the relevance of test samples to unlearned data.
@@
In the updated work, LOKA (Zhang et al., 2025a) introduces multiple storage modules to store distinct knowledge, effectively mitigating conflicts in LLM updating and improving storage efficiency. During training, input knowledge is allocated to the appropriate knowledge memories through similarity-aware knowledge mapping. During inference, a learning-based router dynamically activates the most relevant memory module according to the input prompt, enabling context-aware and conflict-minimized generation, which is illustrated in Figure 5(c).
@@
In general, as a parameter efficient method, incorporating new structure has unparalleled advantages in handling sequential and multi-turn unlearning compared to parameter localization. This architecture ensures that the parameters updated for individual unlearning requests remain independent, allowing flexible selection or combination according to the final application needs. More critically, through parameter integration methods such as fusion mechanisms or learnable routers, it alleviates two crucial problems in continual parameter fine-tuning: catastrophic forgetting of previous knowledge (French, 1999) and knowledge interference between different rounds (Zhang et al., 2025a). However, this plug-in architecture presents several limitations. Firstly, its adaptability to downstream tasks may be constrained. Furthermore, since unlearning is confined solely to the integrated auxiliary structures, deactivating these components can effectively circumvent the defense mechanism, thereby allowing the recovery of unlearned content from the original model (Shen et al., 2025).

### v2

@@ 3.2. Post-Training Unlearning
The main methodology for unlearning in LLMs involves adjusting the parameters of a pre-trained model, a phase typically termed “post-training”. This approach raises two fundamental questions:
(Q1) How should parameters be modified?
(Q2) Which parameters should be targeted for modification?
To address (Q1), following the conventional taxonomy of LLM training, we categorize the unlearning methods into supervised fine-tuning (SFT) and reinforcement learning (RL). Regarding (Q2), the scope of parameter modification may encompass the entire model, a selected subset of parameters, or newly introduced parameters. Strategies for selecting a subset of parameters will be discussed in Section 3.2.3, while methods for integrating new parameters are examined in Section 3.2.4. It is important to note that parameter selection strategies and training algorithms are orthogonal and can be combined flexibly. As a complementary perspective, Section 3.2.5 reviews notable hybrid approaches that integrate multiple techniques discussed in the preceding subsections.
3.2.1. Supervised Fine-tuning (SFT)
Supervised Fine-Tuning (SFT) adapts an LLM to downstream tasks by training it on labeled, task-specific data. The core of an SFT-based unlearning mechanism is the design of its objective for optimization, which dictates how the model’s parameters are adjusted to forget specific knowledge while minimizing the impact on general utility. Different from the unified next-token prediction loss during learning, the design of unlearning objective is quite more diverse. Based on the primary target of the designated objective, we can classify existing methods into three major categories: Text-based, Distribution-based, and Activation-based. Figure 4 compares the general pipeline of different objective categories.
@@ 3.2.1. Supervised Fine-tuning (SFT)
Text-based. These kind of objectives are most intuitive, which aim at minimize or maximize the predicted likelihood of certain text. A representative baseline is Gradient Ascent (jang2022knowledge; Yao2024Large). It reduces the prediction probabilities of forget set samples by directly negating their cross-entropy next-token prediction loss. Some objectives also draw inspiration from preference optimization: NPO (Zhang2024Negative) introduces a reference model to constrain parameter changes, and SimNPO (fan2025simplicity) further improves NPO with length-normalized reward. Another approach focuses on improving likelihood of “substitute responses”, which gives appropriate reply to queries from forget set, without disclosing targeted knowledge (Maini2024TOFU; Mekala2024Alternate). However, these simple baselines often impair general ability of unlearned model, as unlearning corpus typically includes a substantial proportion of tokens that are irrelevant to target knowledge. To alleviate this, some propose to choose only key part in forget set sequences for optimization (Wang2024Selective) or apply different weights to various token positions within the sequences of forget set (Wang2024Rethinking; Feng2024Finegrained). More methods put attention on generating or selecting data. Some utilize an external LLM to generate substitute responses relevant to forget set queries (Mekala2024Alternate; Sinha2024UnStar; xu2025relearn; Liu2024Learning; Gu2024MEOW). Patil2025UPCORE and Chang2025Whicha proposed methods for selecting core subsets from the original unlearning corpus.
@@
Distribution-based. Text-based objectives need to provide labels from limited vocabulary, which restricts optimization space. In order to achieve more fine-grained unlearning, some methods aim to make the model’s output distribution converge to a reference distribution that aligns with the unlearning goals. uses a uniform distribution . Despite using existing distribution like uniform over the entire vocabulary (yuan2024closer), reference distribution can establish either by modifying data or manipulating logits. WHP (Eldan2023Whos) substitutes the unlearning target in original data with unrelated entities to get general knowledge distribution that does not contain sensitive information. WPU (Liu2024Rethinking) improves upon WHP by incorporating diverse substitute entities, performing entity name restoration, and augmenting input prompts. On the other hand, RKLD (Wang2025Balancing) taking the difference between the logits of a model finetuned on the unlearning set and that of original model as reference. Similar logits-aware approach is also adopted by Obliviate (russinovich2025obliviate) and PerMU (wang2025erasing). Distribution-based objectives generally shorten distance to target distribution via minimizing divergence. The most popular choice is KL divergence, but there are also methods using reverse KL (Wang2025Balancing), JS divergence (Singh2025Unlearninga) or f-divergence (Wang2024LLM).
@@
Activation-based. Both text-based and distribution-based methods treat the entire model as a black box, computing losses at the output level and perform back propagation, which is inefficient in many cases. Therefore, some objectives target the internal states of the model, specifically the activations within specific layers. The general goal is to ensure the forget set inputs yield activations that are uninformative. RMU (Li2024WMDP) combines a forget loss that perturbs hidden activations of harmful data towards a fixed random direction. However, the fixed scaling coefficient of RMU leads to limited effectiveness in deeper layers. To overcome this issue,  huu2024effects introduces an adaptive scaling coefficient proportional to the $l_{2}$-norm of the original representation. Similarly, LUNAR (shen2025lunar) aims to redirect the forget data’s activations into a refusal region, so that the model consistently produces safe refusal responses. Guo2024Robust and Wang2024Large take advantage of mechanism interpretability by constructing the activation of the expected answer after unlearning, and reversely calculating the closed-form solution of parameter updates.
@@
Multi-Objective Combination. Beyond unlearning objective itself, most methods add an loss term of retain set in practice to avoid degradation of general performance (Maini2024TOFU; Yao2024Large; Wang2023KGA; yuan2024closer; Lu2024Eraser). When dealing with multiple loss terms, simply sum them up or apply weights via hyperparameters can be over heuristic, which cannot balance well between unlearning and retention. NGDiff (Bu2024Unlearning) treats the combination between objectives as a multi-task optimization problem, achieving a better trade-off through precise normalization and dynamic learning rates. MOLLM (Pan2025MultiObjective) computes a common descent direction in the dual gradient space, yielding an update that simultaneously reduces influence of the target knowledge while preserving overall utility.
3.2.2. Reinforcement Learning (RL)
Reinforcement learning (RL) serves as a foundational methodology in the training of LLMs, enabling them to learn decision-making policies by optimizing the cumulative rewards obtained through environmental interactions. A critical element in applying RL to LLM unlearning involves the design of reward mechanisms tailored to the unlearning scenario.
In pioneering work, Quark (Lu2022QUARK) quantifies undesirable attributes by converting the continuous reward into discrete quantiles and incorporating them as additional reward tokens into the input prompt during RL training. Another approach, DeMem (kassem2023preserving), employs the negative BERTScore (zhang2019bertscore) as a reward signal, which measures the dissimilarity between the model-generated outputs and ground-truth references. Beyond similarity-based rewards, RULE (zhang2025rule) integrates refusal behaviors for data targeted for unlearning into the reward function. Additionally, RULE introduces a rejection steering mechanism for warm-starting the training process, along with a boundary set comprising high-quality hard negatives to sharpen the learning signals near the decision boundary.
As noted in several studies (chu2025sft; chen2025two), RL exhibits a superior generalization capability to unseen data compared to SFT, extending its effectiveness beyond the specific queries encountered during training. A further advantage is that RL-based unlearning operates without altering the underlying loss function, enabling its integration into established RL training pipelines, such as Reinforcement Learning from Human Feedback (RLHF) (bai2022training; ouyang2022training) or RL for enhancing reasoning capabilities (guo2025deepseek). However, a widely acknowledged limitation of RL is its relatively slow convergence rate and inherent training instability (chen2025two).
@@ 3.2.3. Localizing Parameters
     Data based Gradient DEPN (Wu2023DEPN), SURE (zhang2025catastrophic) $\nabla\mathcal{L}_{f}$
SSU (Dou2024Avoiding) ++ gradient of random labeling loss.
Stoehr2024Localizing ++ gradient of KL divergence of output on retain set before/after unlearning.
MemFlex (Tian2024Forget) ++ cosine similarity of $\nabla\mathcal{L}_{f}$ and $\nabla\mathcal{L}_{r}$.
WAGLE (jia2024wagle) ++ element-wise product of $\nabla\mathcal{L}_{f}$ and $\nabla\mathcal{L}_{r}$.
KLUE (yang2025faithun) ++ superficial knowledge regularization.
Activation Selective Pruning (pochinkov2024dissecting) four statistics of activations when processing forget versus retain data.
REVS (Ashuach2024REVS) Combination of activation strength and token association.
FALCON (hu2025falcon) Mutual information of activations of the unlearn and retain set.
     Data free Heuristics RMU (Li2024WMDP) Experimental observation and hyperparameter search optimization.
Adaptive RMU (huu2024effects)
Mechanism LUNAR (shen2025lunar) Knowledge storage mechanism (down-projection matrix of MLP layers) (meng2022locating).
LaW (Wang2024Large)
@@
Simply fine-tuning of all parameters in a model often leads to issues such as high computational costs and potential performance degradation (Dou2024Avoiding). In several studies on interpretability and model editing (meng2022locating; Geva2021Transformer), researchers have demonstrated that knowledge is associated with specific model weights, thus proposing methods to locate relevant parameters for more efficient updates. Various techniques, including causal tracing (meng2022locating), attribution patching (nanda2023attribution), probing (geva2023dissecting), and path patching (goldowsky2023localizing), have been directly applied in research on unlearning (guo2024mechanistic). Furthermore, depending on the specific unlearning scenarios and objectives, numerous studies have proposed different strategies for parameter localization. Based on whether task-specific data are required, we categorize these methods into two distinct classes, as summarized in Table 1.
Data-based.
@@
There are various ways to select certain layers or neurons within the model. The most common selection criterion is the loss gradient w.r.t. the model parameters, calculated on the unlearn set ($D_{u}$), such as approaches in DEPN (Wu2023DEPN) and SURE (zhang2025catastrophic). It is based on the intuition that parameters with larger gradient magnitudes are more influential and should be prioritized for updates. As an updated work, SSU (Dou2024Avoiding) adds a random labeling loss to define a composite loss function, which is a commonly used data augmentation to enhance the stability (neelakantan2015adding). Meanwhile, several works consider the retain set when selecting parameters, reducing the impact of parameter updates on the retain data (Stoehr2024Localizing; Stoehr2024Localizing; Tian2024Forget) (refer to Table 1 for details). Furthermore, yang2025faithun point out that different questions may share the same answer and should avoid unconditionally unlearning the answer regardless of the context. Thus, they propose KLUE, which introduces a superficial knowledge regularization for accurate parameter localization.
@@
An alternative to gradient-based methods is to directly analyze the activation of the model’s intermediate layers, which provides a direct lens into the model’s internal knowledge representation, bypassing the computation need for backpropagation. The method of selective pruning (pochinkov2024dissecting) calculates an importance score for each neuron based on four statistics of its activations when processing unlearn versus retain data. In addition to the activation strength, REVS (Ashuach2024REVS) also considers the rank of a target token when projecting the neuron to the vocabulary space by unembedding matrix, where a lower rank value indicates a stronger association between the target token and the neuron. They show that the combination outperforms methods based solely on activations, token associations, and gradients. Another approach, FALCON (hu2025falcon), uses mutual information of activations of the unlearn and retain set, to identify layers where the hidden representations of forget and retain knowledge are least entangled, targeting these specific layers for modification.
Data-free.
@@
Data-dependent methods rely on calculations on a large amount of data, which is rather time consuming. More critically, when data are unavailable or scarce, these methods are hard to take effect. Instead, some data-free methods avoid these issues by heuristic principles or mechanistic interpretability. Li2024WMDP observe that it is sufficient to compute the loss only on layer $\ell$ and update gradients only on layers $\ell-2$, $\ell-1$ and $\ell$, and perform a hyperparameter search over the layer to select the best layer for fine-tuning. This setting is followed by huu2024effects. Additionally, inspired by insights into knowledge storage mechanism of LLMs (meng2022locating), LUNAR (shen2025lunar) and LaW (Wang2024Large) select the down-projection matrix of the MLP layers to update. In general, heuristic approaches rely on simple and effective rules to select intervention sites, whereas mechanistic approaches target the specific internal circuits responsible for knowledge generation.
@@ 3.2.4. Incorporating New Structure
This type of method generally maintains the original ability of the model by freezing the existing parameters, and achieves forgetting by introducing new parameters or auxiliary structures. A straightforward idea is to insert a new module between two layers of the model and only fine-tune this module, including EUL (Chen2023Unlearn) and GRUN (ren2025general), which is illustrated in Figure 5(a). This module has significantly fewer parameters compared to the original model, sometimes combined with structures like soft gate functions to improve performance (ren2025general). To deal with a sequence of unlearning requests, EUL and GRUN train a separate module on each unlearn task and design a fusion mechanism to merge all modules.
@@
More research focuses on Low-Rank Adaptation (LoRA) (hu2022lora), which adds LoRA adapters to the model (Figure 5(b)). However, standard LoRA lacks sufficient plasticity and often performs poorly in selective unlearning scenarios (cha2024robust), which is followed by several key enhancements. cha2024robust introduce Fisher-weighted Initialization of Low-rank Adapters (FILA). Meanwhile, gao2024large address the challenge of continuous unlearning requests in practical settings. They employ an orthogonal regularization loss to disentangle different unlearning tasks within a single LoRA adapter and additionally train an out-of-distribution (OOD) detector to modulate the adapter activation based on the relevance of test samples to unlearned data.
@@
In the updated work, LOKA (zhang2025resolving) introduces multiple storage modules to store distinct knowledge, effectively mitigating conflicts in LLM updating and improving storage efficiency. During training, input knowledge is allocated to the appropriate knowledge memories through similarity-aware knowledge mapping. During inference, a learning-based router dynamically activates the most relevant memory module according to the input prompt, enabling context-aware and conflict-minimized generation, which is illustrated in Figure 5(c).
@@
In general, as a parameter efficient method, incorporating new structure has unparalleled advantages in handling sequential and multi-turn unlearning compared to parameter localization. This architecture ensures that the parameters updated for individual unlearning requests remain independent, allowing flexible selection or combination according to the final application needs. More critically, through parameter integration methods such as fusion mechanisms or learnable routers, it alleviates two crucial problems in continual parameter fine-tuning: catastrophic forgetting of previous knowledge (french1999catastrophic) and knowledge interference between different rounds (zhang2025resolving). However, this plug-in architecture presents several limitations. Firstly, its adaptability to downstream tasks may be constrained. Furthermore, since unlearning is confined solely to the integrated auxiliary structures, deactivating these components can effectively circumvent the defense mechanism, thereby allowing the recovery of unlearned content from the original model (shen2025lunar).
@@
3.2.5. Composite Approaches
A small portion of post-training unlearning methods integrate multiple techniques discussed in the preceding subsections. We review several notable approaches here as a complementary perspective, categorizing them into (1) parameter arithmetic operations and (2) SAE-based methods.
Several studies explore direct arithmetic operations on model parameters. Inspired by advances in task vectors for knowledge editing (Ilharco2022Editing), some methods fine-tune an intermediate model and combine its parameters arithmetically with those of the original model. For instance, SKU (Liu2024Safer) fine-tunes a “bad model” to obtain a parameter deviation that opposes the unlearning objective. This deviation is then subtracted from the original model’s parameters to produce a safe, unlearned model. A similar strategy is adopted in Eldan2023Whos.
Since training a model of comparable size to the original is computationally intensive, two key refinements have been proposed. First, fine-tuning can be performed using parameter-efficient fine-tuning (PEFT) techniques, where unlearning is achieved by applying negation operations to relevant parameter-efficient modules (PEMs) (zhang2023composing). To further mitigate the risk of degrading general model capabilities, hu2024separate combine an “expert” PEM with an “anti-expert” PEM and derive a general capability vector for preservation. Second, as an alternative to fine-tuning, approximate negative models can be derived via subspace decomposition and projection techniques, such as the Gram–Schmidt orthogonalization used in UNLEARN (Lizzo2024UNLEARN) and the singular value decomposition (SVD) applied in Ethos (Gao2024Ethos).
Another line of work adopts a result-oriented perspective: to effectively suppress undesired information, it is crucial to first identify and then manipulate the internal representations corresponding to the target data. Several studies integrate sparse autoencoders (SAEs) (ng2011sparse) into specific model layers to enhance interpretability and isolate relevant features. For example, farrell2024applying identify features that strongly activate on the unlearn set while minimally impacting the retain set, and then clamp their activations to negative values during inference. Similarly, wu2024codeunlearn introduce a trainable codebook between the encoder and decoder of an SAE. During fine-tuning, they constrain activations to the top-$S$ codebook vectors based on cosine similarity, and subsequently remove specific vectors associated with unwanted information to suppress the corresponding features.
3.3. Inference-Time Unlearning